### v1

# msbm10 scaled1095

Abstract.
While the spectral properties and invariant subspaces of compressed shifts on model spaces are well understood, their behaviour on nearly $S^{*}$-invariant subspaces, a natural generalization with weaker structural constraints, remains largely unexplored. In this paper, we completely characterize the point spectrum, whole spectrum and invariant subspace structure for such compressed shifts by unitary equivalence, using the Frostman shift, Crofoot transform, and Sz.-Nagy–Foias theory. Our results reveal how the relaxation of $S^{*}$-invariance impacts spectral structure and invariant subspaces, bridging a gap between classical model space theory and broader function-theoretic settings.
Key words and phrases:
Nearly invariant subspace, compressed shift, spectrum, truncated Toeplitz operator, Livšic–Moeller theorem, Crofoot transform, functional model
2010 Mathematics Subject Classification:
47A15, 47A10, 30H10.
1. Preliminaries
The compressed shift is a central object in operator theory, arising naturally in the study of Hilbert space contractions. For an inner function $\theta$, the compressed shift $S_{\theta}$ on the model space $K_{\theta}:=H^{2}\ominus\theta H^{2}$ is defined as the compression of the unilateral shift $S\colon f(z)\mapsto zf(z)$ on the Hardy space $H^{2}:=H^{2}(\mathbb{D})$ by
$S_{\theta}:=P_{\theta}S|_{K_{\theta}},$
where $P_{\theta}$ is the orthogonal projection of $L^{2}$ onto the model space $K_{\theta}.$ It is defined by
$(P_{\theta}f)(\lambda)=\langle f,k_{\lambda}^{\theta}\rangle,\;\lambda\in\mathbb{D},f\in L^{2},$
and $k_{\lambda}^{\theta}$ is the reproducing kernel for $K_{\theta},$ expressible in terms of the standard $H^{2}$ kernel $k_{\lambda}(z)$ as
$k_{\lambda}^{\theta}(z)=\frac{1-\overline{\theta(\lambda)}\theta(z)}{1-\overline{\lambda}z}=(1-\overline{\theta(\lambda)}\theta(z))k_{\lambda}(z).$
The operator $S_{\theta}$ highlights the deep relationship between function theory and operator theory, with its structure reflecting the analytic properties of $\theta$. Moreover, it serves as a prototype for all completely non-unitary (c.n.u.) contractions on Hilbert spaces in the Sz.-Nagy–Foias functional model. Specifically, a bounded linear operator $T$ on a Hilbert space $\mathcal{H}$ is defined as a contraction if $\|Tf\|\leq\|f\|$ for all $f\in\mathcal{H}.$ A contraction $T$ is called c.n.u. if there is no nontrivial $T$-reducing subspace $\mathcal{N}\subseteq\mathcal{H}$ such that $T|_{\mathcal{N}}$ is unitary.
The study of invariant subspaces for linear operators has led to profound results, bridging complex function theory, operator theory, and functional analysis. A cornerstone in this field is the Beurling theorem, which characterizes shift-invariant subspaces of the Hardy space $H^{2}$ (see, e.g. [18, A. Corollary 1.4.1]). To extend these ideas to Hardy spaces of an annulus, Hitt introduced the concept of nearly $S^{*}$-invariant subspaces in [14], generalizing Hayashi’s work on the kernels of Toeplitz operators [13]. Further contributions by Sarason [19, 20] deepened the connections with Toeplitz kernels, further enriching the theory.
A closed subspace $\mathcal{M}\subseteq H^{2}$ is called nearly $S^{*}$-invariant (originally weakly invariant) if, for any $f\in\mathcal{M}$ with $f(0)=0$, it follows that $S^{*}f\in\mathcal{M}$. Intuitively, $\mathcal{M}$ is nearly $S^{*}$-invariant if the zeros of its functions can be factored out without leaving the space. Hitt’s seminal work provides a complete characterization of such subspaces in $H^{2}$, later extended to a vectorial setting in [4, Theorem 4.4]. More generally it was extended to the case of nearly $S^{*}$-invariant subspaces with defect $m$ (that is, if $f(0)=0$ then $S^{*}f\in\mathcal{M}+\mathcal{F}$ for a fixed subspace $\mathcal{F}$ of minimal dimension $m$) in [5].
Theorem 1.1.
[14, Proposition 3] The nearly $S^{*}$-invariant subspaces of $H^{2}$ have the form $\mathcal{M}=hK$, with $h\in\mathcal{M}$ of unit norm, $h(0)>0,$ $h$ orthogonal to all elements of $\mathcal{M}$ vanishing at the origin, $K$ an $S^{*}$-invariant subspace, and the operator of multiplication by $h$ isometric from $K$ into $H^{2}$.
The function $h$ is called the extremal function for $\mathcal{M}$ due to its being the unique solution to the extremal problem
$\sup\{\mbox{Re}\;g(0):\;g\in\mathcal{M},\;\|g\|=1\},$
and then either $K$ is trivial or there is an inner function $\theta$ with $\theta(0)=0$ such that $\mathcal{M}=hK_{\theta},$ which is a generalization of $S^{*}$-invariant subspaces.
As shown in [11, Proposition 9.14], the invariant subspaces of $S_{\theta}$ in $K_{\theta}$ are precisely $K_{\theta}\ominus K_{\phi}$, where $\phi$ divides $\theta$. This classical result motivates us to ask a natural question as below.
Question 1: How does this characterization generalize when $K_{\theta}$ is replaced by a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$?
We completely answer Question 1 in Theorem 4.5 and characterize the spectrum of this compressed shift in Theorem 3.9, with both results illustrated in Example 4.6.
To analyze the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, we first revisit the classical compressed shift $S_{\theta}$ on the model space $K_{\theta}$. Recall that $S_{\theta}$, defined as the compression of $S$ to $K_{\theta},$ is the prototypical example of a truncated Toeplitz operator, introduced by Sarason to study function-theoretic properties of model spaces. Given a symbol $g\in L^{2}$, the truncated Toeplitz operator $A_{g}^{\theta}$ is densely defined on bounded $f\in K_{\theta}$ by
$A_{g}^{\theta}f=P_{\theta}(gf),\;\mbox{for}\;f\in\mbox{Dom}(A_{g}^{\theta}),$
with the domain $\mbox{Dom}(A_{g}^{\theta})=\{f\in K_{\theta}:\;gf\in L^{2}\}.$ These operators share many characteristics with classical Toeplitz operators. However, they also exhibit striking differences (see, e.g. [21]). For example, there are bounded truncated Toeplitz operators with unbounded symbols (though any truncated Toeplitz operator with a bounded symbol is itself bounded). Especially, [21, Theorem 3.1] implies $A_{g}^{\theta}=0$ if and only if $g\in\theta H^{2}+\overline{\theta H^{2}}.$
Given a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$, [12, Lemma 2.4] identified the orthogonal projection $P_{\mathcal{M}}$ from $L^{2}$ onto $\mathcal{M}$ as
$P_{\mathcal{M}}f=hP_{\theta}(\overline{h}f),\;\;f\in L^{2}.$
As a corollary, [12, Corollary 2.5] exhibits the reproducing kernel for $\mathcal{M}=hK_{\theta}$ as given by
$k_{\lambda}^{\mathcal{M}}(z):=\overline{h(\lambda)}h(z)\frac{1-\overline{\theta(\lambda)}\theta(z)}{1-\overline{\lambda}z}=\overline{h(\lambda)}h(z)k_{\lambda}^{\theta}(z).$
So, given a measurable function $g$ on $\mathbb{T}$, the truncated Toeplitz operator $A_{g}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ is defined as
$A_{g}^{\mathcal{M}}f=P_{\mathcal{M}}(gf)=hP_{\theta}(\overline{h}gf),\;f\in\mbox{Dom}(A_{g}^{\mathcal{M}}),$
where Dom$(A_{g}^{\mathcal{M}})=\{f\in\mathcal{M}:\;gf\in L^{2}\}.$
The multiplier $M_{h}:K_{\theta}\rightarrow hK_{\theta}$ is a surjective isometry as in Theorem 1.1, and so it is unitary, with adjoint $M_{h}^{*}=M_{h^{-1}},$ and so there follows the result of [12, Theorem 3.1].
Theorem 1.2.
Let $A_{g}^{\mathcal{M}}$ be a bounded operator on $\mathcal{M}=hK_{\theta}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$. Then there is a unitary equivalence between $A^{\mathcal{M}}_{g}$ and $A^{\theta}_{|h|^{2}g}$ given by $A_{g}^{\mathcal{M}}=M_{h}A^{\theta}_{|h|^{2}g}M_{h}^{*}.$
Proof.
Taking $f\in\mathcal{M}=hK_{\theta}$, it holds that
$A^{\mathcal{M}}_{g}f=hP_{\theta}(\overline{h}gf)=M_{h}A^{\theta}_{|h|^{2}g}(h^{-1}f)=M_{h}A_{|h|^{2}g}^{\theta}M_{h}^{*}f,$
ending the proof.∎
We note that Theorem 1.2 and [21, Theorem 3.1] provide a characterization of zero truncated Toeplitz operators on $\mathcal{M}=hK_{\theta}$.
Corollary 1.3.
Let $\mathcal{M}=hK_{\theta}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$ and $|h|^{2}g\in L^{2},$ then $A_{g}^{\mathcal{M}}=0$ on $\mathcal{M}$ if and only if $|h|^{2}g$ belongs to $\theta H^{2}+\overline{\theta}\overline{H^{2}}.$
Kernel spaces are crucial for understanding the structure of linear operators, especially Toeplitz operators and truncated Toeplitz operators (see, e.g. [1, 21]). [16, Corollary 3.2] (with $n=1$) states that the kernel of a truncated Toeplitz operator on model space is nearly $S^{*}$-invariant with defect at most $1$. This result can be extended to the operator $A^{\mathcal{M}}_{g}$ acting on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
Corollary 1.4.
Let $A^{\mathcal{M}}_{g}$ be a bounded truncated Toeplitz operator on a nearly $S^{*}$-invariant space $\mathcal{M}=hK_{\theta}\neq\{0\}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$. Then $\ker A^{\mathcal{M}}_{g}$ is nearly $S^{*}$-invariant with defect at most $1$.
Proof.
Suppose that $f\in\ker A^{\mathcal{M}}_{g}$ and $f(0)=0$. We have $f=hk$ with some $k\in K_{\theta}$ and we note that by near $S^{*}$-invariance, it follows that $h(0)\neq 0$. Otherwise every function in $\mathcal{M}$ would vanish at 00, which yields =$\mathcal{M}=\{0\}$. So it follows that $k(0)=0.$ From $A^{\mathcal{M}}_{g}f=0$, Theorem 1.2 implies
$M_{h}A_{|h|^{2}g}^{\theta}(M_{h}^{*}f)=M_{h}A_{|h|^{2}g}^{\theta}(h^{-1}f)=M_{h}A_{|h|^{2}g}^{\theta}(k)=0$
and so $k\in\ker A^{\theta}_{|h|^{2}g}$. Hence [16, Corollary 3.2] asserts that there is a one-dimensional subspace $\mathcal{F}=\langle x\rangle$, independent of $k$, such that $S^{*}k=\overline{z}k\in\ker A^{\theta}_{|h|^{2}g}+\mathcal{F}$. Then there is a $\lambda\in\mathbb{C}$ such that $A^{\theta}_{|h|^{2}g}(\overline{z}k-\lambda x)=0.$ We can express this as
$h^{-1}A_{g}^{\mathcal{M}}h(\overline{z}k-\lambda x)=h^{-1}A_{g}^{\mathcal{M}}(\overline{z}f-\lambda hx)=0.$
That is, $S^{*}f-\lambda(hx)\in\ker A^{\mathcal{M}}_{g}$, as required. ∎
Finally, we demonstrate that the truncated Toeplitz operator $A_{g}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ are $D$-symmetric with respect to the conjugation $D:=M_{h}CM_{h}^{*}$, where
$Cf=\theta\overline{zf},$ (1.1)
is a bijective map from $\theta H^{2}$ to $\overline{zH^{2}}$ which also takes $K_{\theta}$ to itself.
Remark 1.5.
Let $A_{g}^{\mathcal{M}}$ be a bounded operator on $\mathcal{M}=hK_{\theta}$ with extremal function $h$ and associated inner function $\theta$, $\theta(0)=0$, then $DA_{g}^{\mathcal{M}}D=A_{\overline{g}}^{\mathcal{M}}$ with $D=M_{h}CM_{h}^{*}.$
Proof.
Theorem 1.2 implies $M_{h}^{*}A_{g}^{\mathcal{M}}M_{h}=A^{\theta}_{|h|^{2}g},$ which further yields that
$CM_{h}^{*}A_{g}^{\mathcal{M}}M_{h}C=CA^{\theta}_{|h|^{2}g}C=A^{\theta}_{|h|^{2}\overline{g}}.$
Note $M_{h}^{*}A_{\overline{g}}^{\mathcal{M}}M_{h}=A^{\theta}_{|h|^{2}\overline{g}},$ and so
$CM_{h}^{*}A_{g}^{\mathcal{M}}M_{h}C=CA^{\theta}_{|h|^{2}g}C=M_{h}^{*}A_{\overline{g}}^{\mathcal{M}}M_{h}.$
That is to say, $DA_{g}^{\mathcal{M}}D=(M_{h}CM_{h}^{*})A_{g}^{\mathcal{M}}(M_{h}CM_{h}^{*})=A_{\overline{g}}^{\mathcal{M}}.$
Building on the above foundational properties of $A_{g}^{\mathcal{M}}$ on $\mathcal{M}=hK_{\theta}$, we now turn to address Question 1. In the specific case where $g(z)=z,$ Theorem 1.2 establishes that the compressed shift $A_{z}^{\mathcal{M}}$ on $\mathcal{M}=hK_{\theta}$ satisfies the following key property
$A_{z}^{\mathcal{M}}=M_{h}A_{|h|^{2}z}^{\theta}M_{h}^{*}.$ (1.2)
Let $(A_{z}^{\mathcal{M}})^{*}$ denote the adjoint of $A_{z}^{\mathcal{M}}$, then for $f,g\in\mathcal{M},$ it holds that
$\langle(A_{z}^{\mathcal{M}})^{*}f,g\rangle=\langle f,A_{z}^{\mathcal{M}}g\rangle=\langle f,P_{\mathcal{M}}(Sg)\rangle=\langle f,Sg\rangle=\langle P_{\mathcal{M}}(\overline{z}f),g\rangle,$
yielding that $(A_{z}^{\mathcal{M}})^{*}=A_{\overline{z}}^{\mathcal{M}}.$ Meanwhile, it holds that
$(A_{z}^{\mathcal{M}})^{*}=M_{h}A_{|h|^{2}\overline{z}}^{\theta}M_{h}^{*}.$ (1.3)
Equations (1.2) and (1.3) reveal that the spectral properties and invariant subspaces of the compressed shift $A_{z}^{\mathcal{M}}$ (respectively, its adjoint $(A_{z}^{\mathcal{M}})^{*}$) on $\mathcal{M}=hK_{\theta}$ are fully determined by those of the truncated Toeplitz operator $A_{|h|^{2}z}^{\theta}$ (respectively, $A_{|h|^{2}\bar{z}}^{\theta}$) on $K_{\theta}$, where $\theta(0)=0$. Specifically, the spectrum of $A_{z}^{\mathcal{M}}$ (respectively, $(A_{z}^{\mathcal{M}})^{*}$) coincides with that of $A_{|h|^{2}z}^{\theta}$ (respectively, $A_{|h|^{2}\bar{z}}^{\theta}$) i.e. $\sigma(A_{z}^{\mathcal{M}})=\sigma(A_{|h|^{2}z}^{\theta})$ (respectively, $\sigma(A_{\overline{z}}^{\mathcal{M}})=\sigma(A_{|h|^{2}\overline{z}}^{\theta})$). Meanwhile, every invariant subspace of $A_{z}^{\mathcal{M}}$ (respectively, $(A_{z}^{\mathcal{M}})^{*}$) is of the form $M_{h}(\mathcal{N})$, where $\mathcal{N}\subseteq K_{\theta}$ is an invariant subspace of $A_{|h|^{2}z}^{\theta}$ (respectively, $A_{|h|^{2}\bar{z}}^{\theta}$ ).
This paper is structured as follows: First, we derive explicit representations for the truncated Toeplitz operator $A_{|h|^{2}z}^{\theta}$ and its adjoint $A_{|h|^{2}\bar{z}}^{\theta}$ on the model space $K_{\theta}$ with $\theta(0)=0$ in Propositions 2.3 and 2.1, establishing fundamental results on their cyclic vectors and actions on kernel functions in Propositions 2.4 and 2.5. Next, we apply unitary equivalence, using the Frostman shift, Crofoot transform, and Sz.-Nagy–Foias theory to rigorously analyze the spectral properties of $A_{z}^{\mathcal{M}}$, identifying its essential spectrum in Corollary 3.2, point spectrum in Theorem 3.8 and whole spectrum in Theorem 3.9. Finally, we focus on finding the generalized eigenvectors for any eigenvalue of $A_{|h|^{2}\bar{z}}^{\theta}$ with multiplicity $n\geq 2$ to display more finite-dimensional invariant subspaces in Proposition 4.3, and then classify all invariant subspaces for the compressed shift $A_{z}^{\mathcal{M}}$ acting on an arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$ in Theorem 4.5. Especially, we demonstrate Theorems 3.9 and 4.5 in Example 4.6, thereby unifying these results within a generalized framework of model space theory.
2. The compressed shifts on nearly invariant subspaces
Given an inner function $\theta$ with $\theta(0)=0,$ there are two orthogonal direct sums
$K_{\theta}=K_{z}\oplus zK_{\theta\overline{z}}=K_{\theta\overline{z}}\oplus\mathbb{C}\theta\overline{z},$ (2.1)
and the backward shift $S^{*}$ acts as $0\oplus M_{\overline{z}}$ with respect to the first sum. Next
$\langle A_{|h|^{2}\bar{z}}^{\theta}f,g\rangle=\langle|h|^{2}\overline{z}f,g\rangle=\langle h\overline{z}f,hg\rangle=\langle\bar{z}f,g\rangle$
if $\overline{z}f\in K_{\theta},$ i.e., if $f\in zK_{\theta\bar{z}}$ and $g\in K_{\theta}$ is arbitrary. That is, on the subspace $zK_{\theta\bar{z}}$ these two operators coincide. Then we need to work out $A_{|h|^{2}\bar{z}}^{\theta}1$ to understand $A_{|h|^{2}\bar{z}}^{\theta}$ on $K_{\theta}.$ Indeed,
$\langle A_{|h|^{2}\bar{z}}^{\theta}1,g\rangle=\langle|h|^{2}\bar{z}1,g\rangle=\langle h1,hgz\rangle=\langle 1,gz\rangle=0$ (2.2)
if $gz\in K_{\theta}.$ The second decomposition in (2.1) together with (2.2) show that $A_{|h|^{2}\bar{z}}^{\theta}1\in\mathbb{C}\theta{\bar{z}}$. That is, $A_{|h|^{2}\bar{z}}^{\theta}1$ is a multiple of $\theta\bar{z}$. Considering $\langle|h|^{2}\bar{z}1,\theta\bar{z}\rangle=\langle|h|^{2},\theta\rangle,$ we finally conclude that
$A_{|h|^{2}\bar{z}}^{\theta}1=\bar{z}\theta\langle|h|^{2},\theta\rangle.$ (2.3)
So we have established the following proposition.
Proposition 2.1.
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Then
$B_{v}f=\left\{\begin{array}[]{ll}\bar{z}f,&f\in zK_{\theta\overline{z}},\\ \overline{v}\theta\overline{z},&f=1.\end{array}\right.$ (2.6)
Under the first decomposition in (2.1), (2.6) yields $B_{v}$ as a rank-1 perturbation of $S_{\theta}^{*}$ on $K_{\theta}.$ Moreover, the constraints of the extremal function $h\in\mathcal{M}=hK_{\theta}$ ensure that $|v|\leq 1$.
Here we include an example with inner $\theta$ having repeated zeros to exhibit the difference between $B_{v}$ and $S_{\theta}^{*}$ acting on $K_{\theta}.$ Our convention is that the $j$-th column in the matrix contains the image of the $j$-th vector, so if $S^{*}v_{j}=\sum_{k=1}^{n}c_{k}v_{k},$ then the coefficients $c_{1},\cdots,c_{n}$ are put in the $j$-th column.
Example 2.2.
Let $\theta(z)=z^{4}$ ($z^{n}$ is similar). With respect to the orthonormal basis $\{1,z,z^{2},z^{3}\}$ of $K_{\theta}$ we have the matrix
$[S_{\theta}^{*}]=\left(\begin{array}[]{cccc}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ 0&0&0&0\end{array}\right).$
The only eigenvalue is 00 and the invariant subspaces are model subspaces of $K_{\theta}$, i.e., $\{0\},K_{z},K_{z^{2}},K_{z^{3}}$ and $K_{z^{4}}$. On the other hand, the matrix of $B_{v}$ is
$[B_{v}]=\left(\begin{array}[]{cccc}0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \alpha&0&0&0\end{array}\right)$
where $\alpha=\langle|h|^{2},z^{4}\rangle$. Note that $[B_{v}]^{4}=\alpha I$ and there are four distinct eigenvalues, the fourth roots of $\alpha$ (in the case $\alpha\neq 0$). So the lattice of invariant subspaces is rather different, being the linear spans of eigenvectors. Note that the two lattices of invariant subspaces are very different, even though the matrices differ in only one entry.
We now analyse the action of $A_{|h|^{2}z}^{\theta}$ on $K_{\theta}$ when $\theta(0)=0.$ Considering $M_{h}:K_{\theta}\rightarrow hK_{\theta}$ is unitary in Theorem 1.1, we obtain
$\langle hf,hg\rangle=\langle f,g\rangle,\;\mbox{for}\;f,g\in K_{\theta}.$
Next, we deduce
$\langle A_{|h|^{2}z}^{\theta}f,g\rangle=\langle hzf,hg\rangle,\;\mbox{for}\;f,g\in K_{\theta}.$
Thus, if $zf\in K_{\theta}$ we conclude that $A_{|h|^{2}z}^{\theta}f=zf$ by taking $g$ as the reproducing kernel. Employing the second orthogonal decomposition in (2.1), it follows that $zf\in K_{\theta}$ for all $f\in K_{\theta\bar{z}}.$ That means $A_{|h|^{2}z}^{\theta}$ and $S$ coincide on the subspace $K_{\theta\bar{z}}.$
After that we use conjugation $C$ in (1.1) to compute $A_{|h|^{2}z}^{\theta}(\theta\bar{z})$. Since $1\in K_{\theta}$ due to $\theta(0)=0,$ $C1=\theta\bar{z}$, (2.3) yields that
$A_{|h|^{2}z}^{\theta}(\theta\bar{z})$ $=A_{|h|^{2}z}^{\theta}C1=CB_{v}1=C(\bar{z}\theta\langle|h|^{2},\theta\rangle)=\theta\overline{z\theta\langle|h|^{2},\theta\rangle}z=v.$
Hence we summarize the above calculations in the following proposition.
Proposition 2.3.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Then
$A_{v}f=\left\{\begin{array}[]{ll}zf,&f\in K_{\theta\overline{z}},\\ v,&f\in\theta\overline{z}.\end{array}\right.$ (2.9)
Equation (2.9) indicates how $A_{v}$ is a rank-1 perturbation of $S_{\theta}$ on $K_{\theta}.$ For $S_{\theta}$ on $K_{\theta}$, [11, Proposition 9.13] shows that it is a cyclic operator with cyclic vector $k_{0}=1-\overline{\theta(0)}\theta.$ The next proposition focuses on the cyclic vectors of $A_{v}$ and $B_{v}$.
Proposition 2.4.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. Then both $A_{v}$ and its adjoint $B_{v}$ are cyclic operators on $K_{\theta}$, with cyclic vectors $1$ and $\theta/z$.
Proof.
Let $u=\theta/z$. For the operator $B_{v},$ we observe that
$B_{v}u=S^{*2}\theta+u(0)\overline{v}u=S^{*}u+u(0)\overline{v}u,$
which shows $\bigvee\{u,B_{v}u\}$ contains $S^{*}u$. By induction, $\bigvee\{B_{v}^{n}u:n\geq 0\}$ contains $\bigvee\{S^{*n}u:n\geq 0\}$. Since $\bigvee\{S^{*n}\theta:\;n\geq 1\}=K_{\theta}$ (see, e.g. [11, Proposition 5.15]) and $\theta(0)=0,$ we conclude that
$\bigvee\{B_{v}^{n}u:\;n\geq 0\}=\bigvee\{S^{*n}u:\;n\geq 0\}=K_{\theta}.$ (2.10)
Similarly, we obtain
$\bigvee\{B_{v}^{n}1:\;n\geq 0\}=\bigvee\{1,S^{*n}u:\;n\geq 0\}=K_{\theta}.$ (2.11)
For the operator $A_{v}=CB_{v}C$, where $C$ is the conjugation in (1.1), we have
$\bigvee\{A_{v}^{n}1:\;n\geq 0\}$ =$=\bigvee\{(CB_{v}C)^{n}1:\;n\geq 0\}$
=$=\bigvee\{CB_{v}^{n}C1:\;n\geq 0\}$
$=C\bigvee\{B_{v}^{n}u:\;n\geq 0\}.$
This, together with (2.10), implies $\bigvee\{A_{v}^{n}1:\;n\geq 0\}=CK_{\theta}=K_{\theta}.$
Likewise, by (2.11) it yields that
$\bigvee\{A_{v}^{n}u:\;n\geq 0\}$ =$=\bigvee\{(CB_{v}C)^{n}u:\;n\geq 0\}$
=$=\bigvee\{CB_{v}^{n}Cu:\;n\geq 0\}$
$=C\bigvee\{B_{v}^{n}1:\;n\geq 0\}$
$=C\bigvee\{1,\;B_{v}^{n}u:\;n\geq 0\}=K_{\theta}.$
Thus both $1$ and $\theta/z$ are cyclic vectors for $A_{v}$ and $B_{v}$ on $K_{\theta}$ with $\theta(0)=0.$
A detailed analysis of the operators’ actions on kernel functions is conducted to further understand their properties. We will use the notation
$\widetilde{k_{\lambda}^{\theta}}=Ck_{\lambda}^{\theta}=\frac{\theta(z)-\theta(\lambda)}{z-\lambda}$ (2.12)
with the conjugation $C$ given in (1.1).
Proposition 2.5.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ and $B_{v}=A_{v}^{*}$.
$(a)$ For $w\in\mathbb{D}$,
$B_{v}k_{w}^{\theta}=\theta\bar{z}\overline{v}+\overline{w}k_{w}^{\theta/z},\;\;A_{v}\widetilde{k_{w}^{\theta}}=v+wz\widetilde{k_{w}^{\theta/z}}.$
$(b)$ For $w\in\mathbb{D}\setminus\{0\},$
$A_{v}k_{w}^{\theta}=v\frac{\overline{\theta(w)}}{\overline{w}}+zk_{w}^{\theta/z},\;\;B_{v}\widetilde{k_{w}^{\theta}}=\frac{\theta(w)}{w}\overline{v}\theta\overline{z}+\widetilde{k_{w}^{\theta/z}}.$
Proof.
Let $k_{w}^{\theta}$ denote the reproducing kernel for $K_{\theta}$ at the point $w\in\mathbb{D}$. Now
$k_{w}^{\theta}=\langle k_{w}^{\theta},1\rangle+k_{w}^{\theta}-\langle k_{w}^{\theta},1\rangle 1=1+(k_{w}^{\theta}-1)$
where the second term is in the orthogonal complement of $K_{z}$ in $K_{\theta},$ that is, $zK_{\theta/z}.$ Thus
$B_{v}k_{w}^{\theta}=\theta\bar{z}\overline{v}+\bar{z}(k_{w}^{\theta}-1),$
which together with
$\bar{z}(k_{w}^{\theta}-1)=\overline{w}k_{w}^{\theta/z}$ (2.13)
imply the expression of $B_{v}k_{w}^{\theta}.$
Now let us look at $A_{v}$, using the fact $(A_{v})^{*}=B_{v}$, for $\mu,w\in\mathbb{D}$ we have
$\langle A_{v}k_{w}^{\theta},k_{\mu}^{\theta}\rangle$ $=\langle k_{w}^{\theta},B_{v}k_{\mu}^{\theta}\rangle=\langle k_{w}^{\theta},\bar{z}\theta\overline{v}+\bar{z}(k_{\mu}^{\theta}-1)\rangle$
$=v\frac{\overline{\theta(w)}}{\overline{w}}+\frac{\overline{k_{\mu}^{\theta}(w)}-1}{\overline{w}}$
$=v\frac{\overline{\theta(w)}}{\overline{w}}+\frac{k_{w}^{\theta}(\mu)-1}{\overline{w}}$
and by (2.13) we deduce that
$A_{v}k_{w}^{\theta}=v\frac{\overline{\theta(w)}}{\overline{w}}+\frac{k_{w}^{\theta}-1}{\overline{w}}=v\frac{\overline{\theta(w)}}{\overline{w}}+zk_{w}^{\theta/z}.$
By the conjugation $C$ in (1.1), we deduce that
$A_{v}\widetilde{k_{w}^{\theta}}=CB_{v}k_{w}^{\theta}=v+wz\widetilde{k_{w}^{\theta/z}}.$
It further entails that
$B_{v}\widetilde{k_{w}^{\theta}}=CA_{v}k_{w}^{\theta}=\frac{\theta(w)}{w}\overline{v}\theta\bar{z}+\widetilde{k_{w}^{\theta/z}}.$
3. Spectra of the compressed shifts on nearly invariant subspaces
In this section, we concentrate on determining the spectrum of the compressed shift $A_{z}^{\mathcal{M}}$ on an arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is an associated inner function satisfying $\theta(0)=0$. By virtue of (1.2), this problem reduces to characterizing the spectrum of the operator $A_{v}$ on the model space $K_{\theta}$ with $\theta(0)=0$. For an inner function $\theta\in H^{2}$, it can be expressed as
[…]
where $y\in[0,2\pi)$, the factor $B$ is a Blaschke product with zero set $\{a_{n}\}_{n\geq 1}$ (counting multiplicities), and $s_{\mu}$ is a singular inner function with corresponding positive singular measure $\mu$. We now define the spectrum $\sigma(\theta)$ in terms of the above canonical factorization of $\theta$.
Definition 3.1.
For a nonconstant inner function $\theta=Bs_{\mu}$, the spectrum of $\theta$ is the set
$\sigma(\theta)=\{a_{n}\}_{n\geq 1}^{-}\cup\operatorname{supp}\mu.$
[…]
$\sigma(\theta)=\{\lambda\in\overline{\mathbb{D}}:\;\lim_{z\in\mathbb{D}}\inf_{z\rightarrow\lambda}|\theta(z)|=0\}.$
For the compressed shift $S_{\theta}$ on a model space $K_{\theta}$, Livšic and Moeller proved that $\sigma(S_{\theta})=\sigma(\theta)$ (see [15, 17]). The spectrum $\sigma(T)$ of a linear and bounded operator $T$ is defined as
$\sigma(T)=\{\lambda\in\mathbb{C}:\;T-\lambda I\;\mbox{is not invertible}\}.$
In addition, a linear bounded operator $T$ is said to be Fredholm if its range $R(T)$ is closed and both $R(T)^{\perp}$ and the null space $N(T)$ are finite-dimensional. The essential spectrum is defined as $\sigma_{e}(T)=\{\lambda\in\mathbb{C}:\;T-\lambda I\;\mbox{is not Fredholm}\}$, which is invariant under perturbation by compact operators (see [22]). In the case of $S_{\theta}$, its essential spectrum is given by $\sigma_{e}(S_{\theta})=\sigma(\theta)\cap\mathbb{T}$, where $\mathbb{T}$ denotes the unit circle (see e.g. [11, Proposition 9.26] or [21, Lemma 2.5]). Then Proposition 2.3 implies $\sigma_{e}(A_{v})=\sigma(\theta)\cap\mathbb{T}.$ This result, combined with (1.2), leads to the following corollary on the essential spectrum of $A_{z}^{\mathcal{M}}$.
Corollary 3.2.
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0$; then $\sigma_{e}(A_{z}^{\mathcal{M}})=\sigma(\theta)\cap\mathbb{T}.$
Next, the eigenvalues together with eigenvectors and then the whole spectrum are explored for $A_{v}$ in (2.9) on $K_{\theta}$ with $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$. For the case $v=0$, $A_{v}=S_{\theta}$ on $K_{\theta}.$ And for the case $|v|=1$, $A_{v}$ is the same as the unitary operator $U_{v}$ in Clark’s paper [7] with $\theta(0)=0.$ Since the $v=0$ case is well-understood and our subsequent proofs remain valid for this case, we shall henceforth assume $|v|<1$ without loss of generality. At this time, $A_{v}$ is a c.n.u. contraction on a model space $K_{\theta}$ with $\theta(0)=0.$ In consequence, we apply the Sz.-Nagy–Foias theory of characteristic functions for contractions on Hilbert space to delve our deeper exploration.
Recall from the book [23] that the characteristic function of a contraction $T$ in a Hilbert space $\mathcal{H}$ is a triple $(\mathfrak{D}_{T},\mathfrak{D}_{T^{*}},\Theta_{T}(\lambda))$ where
$D_{T}$ $=\sqrt{I-T^{*}T},\quad D_{T^{*}}=\sqrt{I-TT^{*}},$
$\mathfrak{D}_{T}$ $=\overline{D_{T}\mathcal{H}},\quad\mathfrak{D}_{T^{*}}=\overline{D_{T^{*}}\mathcal{H}}\quad\text{and}$
$\Theta_{T}(\lambda)$ $=\left[-T+\lambda D_{T^{*}}(1-\lambda T^{*})^{-1}D_{T}\right]\bigg{|}\mathfrak{D}_{T}.$
$\mathfrak{D}_{T}$ and $\mathfrak{D}_{T^{*}}$ are the defect spaces of $T$. Let us first proceed with the concrete analysis on the defect spaces of $A_{v}$ with $|v|<1$.
Lemma 3.3.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$, $B_{v}=A_{v}^{*}$. Then $\mathfrak{D}_{A_{v}}=\mathbb{C}\theta/z,\;\mathfrak{D}_{B_{v}}=\mathbb{C}$.
Proof.
For any $f\in K_{\theta},$ there is the decomposition $f=(f-f(0))+f(0)\in zK_{\theta\overline{z}}\oplus K_{z}$. Combining Propositions 2.3 and 2.1, we deduce that
$A_{v}B_{v}f$ $=A_{v}B_{v}(f-f(0)+f(0))$
$=(f-f(0))+|v|^{2}f(0),$
which yields that
$(I-A_{v}B_{v})f=(1-|v|^{2})f(0)=(1-|v|^{2})k_{0}^{\theta}\otimes k_{0}^{\theta}(f)=(1-|v|^{2})P_{K_{z}}(f).$ (3.1)
Thus we conclude that
$\sqrt{I-A_{v}B_{v}}=\sqrt{1-|v|^{2}}P_{K_{z}}.$
Furthermore, using $CA_{v}C=B_{v}$ and (3.1), it yields that
$(I-CA_{v}CCB_{v}C)f$ $=(1-|v|^{2})Ck_{0}^{\theta}\otimes k_{0}^{\theta}Cf$
$=(1-|v|^{2})\langle Cf,k_{0}^{\theta}\rangle Ck_{0}^{\theta}$
$=(1-|v|^{2})\langle f,\theta/z\rangle\theta/z$
$=(1-|v|^{2})P_{\theta/zK_{z}}f,$
which implies that
$\sqrt{I-B_{v}A_{v}}=\sqrt{1-|v|^{2}}P_{\theta/zK_{z}}.$
Hence the two defect spaces are $\mathfrak{D}_{A_{v}}=\mathbb{C}\theta/z$ and $\mathfrak{D}_{B_{v}}=\mathbb{C}$ with $\mathop{\rm dim}\nolimits\mathfrak{D}_{A_{v}}=\mathop{\rm dim}\nolimits\mathfrak{D}_{B_{v}}=1.$
Considering [23, Theorem 2.3] in the context of our c.n.u. operator $A_{v}$ with $|v|<1$, we observe that $A_{v}$ is a special case of $Z(A)$ in [10, Definition 2.1] with $A\xi=v\xi$ for $\xi\in N=\mathbb{C}$. And then we may use [10, Theorem 3.2] and Lemma 3.3 to deduce that $A_{v}$ ($|v|<1$) on $K_{\theta}$ with $\theta(0)=0$ has the characteristic function
$\left\{\mathbb{C}\frac{\theta}{z},\;\mathbb{C},\;\Theta_{A_{v}}(\lambda)\right\}$
where $\Theta_{A_{v}}(\lambda)$ is an analytic operator-valued function whose values are contractive operators from $\mathfrak{D}_{A_{v}}=\mathbb{C}\frac{\theta}{z}$ into $\mathfrak{D}_{A_{v}^{*}}=\mathbb{C}$ and its definition is
$\Theta_{A_{v}}(\lambda)\left(\xi\frac{\theta}{z}\right)=\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}\xi,\;~{}\xi\in\mathbb{C}.$ (3.2)
Applying [23, Theorem 2.3] to our operator $A_{v}$ with $|v|<1$, it yields that $A_{v}$ is unitarily equivalent to the canonical compressed shift $S$ on the model space $H^{2}(\mathbb{C})\ominus\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)$, where $H^{2}(\mathbb{C})=H^{2}(\mathbb{D}).$ From (3.2), we can write
$\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)$ =$=\left\{\Theta_{A_{v}}(\lambda)\left(\frac{\theta}{z}f(z)\right):\;f\in H^{2}(\mathbb{C})\right\}$
=$=\left\{\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}f(z):\;f\in H^{2}(\mathbb{C})\right\}$
$=\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}H^{2}(\mathbb{C}).$
So it follows that
$H^{2}(\mathbb{C})\ominus\Theta_{A_{v}}(\lambda)H^{2}(\mathbb{C}\theta/z)=H^{2}(\mathbb{C})\ominus\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}H^{2}(\mathbb{C})=K_{\theta_{v}}$
with $\theta_{v}$ is a Frostman shift of the inner function $\theta$ defined by
$\theta_{v}(\lambda)=\frac{\theta(\lambda)-v}{1-\overline{v}\theta(\lambda)}.$ (3.3)
Therefore, we obtain the unitary model for $A_{v}$ with $|v|<1$ on $K_{\theta}$ with $\theta(0)=0$ as below.
Lemma 3.4.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta,$ $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Then $A_{v}$ is unitarily equivalent to the compressed shift $S_{\theta_{v}}$ on a model space $K_{\theta_{v}}$, where $\theta_{v}$ is a Frostman shift of $\theta$.
It is shown in [8] (see also [6, Proposition 4.3] or [21, Theorem 13.1]) that for an inner function $\theta$ and $v\in\mathbb{D}$, the Crofoot transform
$J_{v}f=\frac{\sqrt{1-|v|^{2}}}{1-\overline{v}\theta}f$
defines a unitary operator from the model space $K_{\theta}$ onto the model space $K_{\theta_{v}}$ with $\theta_{v}$ defined in (3.3). Moreover, its inverse of $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is
$J_{v}^{-1}g=\frac{\sqrt{1-|v|^{2}}}{1+\overline{v}\theta_{v}}g,\;\mbox{for}\;g\in K_{\theta_{v}}.$ (3.4)
By the unitary mapping $J_{v}:\;K_{\theta}\rightarrow K_{\theta_{v}},$ the following commutative diagram is derived in [21, Lemma 13.3]: that is, $J_{v}A_{v}=S_{\theta_{v}}J_{v}$.
[…]
For the compressed shift $S_{\theta}$ on $K_{\theta}$, [21, Lemma 2.5] or [11, Corollary 9.24] exhibits its point spectrum as
$\sigma_{p}(S_{\theta})=\sigma(\theta)\cap\mathbb{D}=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=0\}.$ (3.6)
Meanwhile, the corresponding eigenvectors are found in [11, Corollary 9.25] by showing
$\ker(S_{\theta}-\lambda I)=\mathbb{C}Ck_{\lambda}^{\theta},\;\ker(S_{\theta}^{*}-\overline{\lambda}I)=\mathbb{C}k_{\lambda}^{\theta}$ (3.7)
with conjugation $C$ in (1.1) and $\lambda\in\mathbb{D}$ satisfying $\theta(\lambda)=0.$ So we can use them to describe the point spectrum and corresponding eigenvectors of $A_{v}$ with $|v|<1$ on $K_{\theta}$ with $\theta(0)=0$ by (3.5).
Proposition 3.5.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Then the point spectrum is given by
$\sigma_{p}(A_{v})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}.$
For each $\lambda\in\sigma_{p}(A_{v})$, the corresponding eigenvector is
$Ck_{\lambda}^{\theta}=(\theta(z)-v)/(z-\lambda),$
where $C$ is the conjugation on $K_{\theta}.$
Proof.
By (3.5), we have the unitary equivalence
$S_{\theta_{v}}=J_{v}A_{v}J_{v}^{-1},$ (3.8)
where $S_{\theta_{v}}$ denotes the compressed shift on $K_{\theta_{v}}.$ This immediately yields $\sigma_{p}(A_{v})=\sigma_{p}(S_{\theta_{v}}).$
Applying (3.6) to the Frostman shift $\theta_{v}$ defined in (3.3), we obtain
$\sigma_{p}(S_{\theta_{v}})=\sigma(\theta_{v})\cap\mathbb{D}=\{\lambda\in\mathbb{D}:\;\theta_{v}(\lambda)=0\}=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}.$
This has also been established in [21, Theorem 14.1].
For the eigenvector, observe that by (3.7), we have $S_{\theta_{v}}(Ck_{\lambda}^{\theta_{v}})=\lambda(Ck_{\lambda}^{\theta_{v}}).$ Combining this with the unitary equivalence (3.8) yields
$A_{v}J_{v}^{-1}(Ck_{\lambda}^{\theta_{v}})=\lambda J_{v}^{-1}(Ck_{\lambda}^{\theta_{v}}).$ (3.9)
Through direct calculations using (3.4), (3.3) and (2.12), we derive that
$J_{v}^{-1}(Ck_{\lambda}^{\theta_{v}})$ $=J_{v}^{-1}\left(\frac{\theta_{v}(z)-0}{z-\lambda}\right)=\frac{\sqrt{1-|v|^{2}}}{1+\overline{v}\theta_{v}}\frac{\theta_{v}(z)}{z-\lambda}$
$=\frac{1}{\sqrt{1-|v|^{2}}}\frac{\theta-v}{z-\lambda}=\frac{1}{\sqrt{1-|v|^{2}}}Ck_{\lambda}^{\theta}.$ (3.10)
Substituting (3.10) into (3.9), we complete the proof. ∎
Applying $B_{v}=(A_{v})^{*}$ and $B_{v}=CA_{v}C$, there follows a remark on $B_{v}$’s point spectrum.
Remark 3.6.
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Then the point spectrum is given by
$\sigma_{p}(B_{v})=\{\overline{\lambda}\in\mathbb{D}:\;\theta(\lambda)=v\}.$
For each $\overline{\lambda}\in\sigma_{p}(B_{v})$, the corresponding eigenvector is $k_{\lambda}^{\theta}.$
Two illustrative examples demonstrate the eigenvalues and eigenvectors of $A_{v}$ and $B_{v}$.
Example 3.7.
(1)  Let $\theta(z)=z^{2}\frac{z-a}{1-\overline{a}z}$ when $10<|a|<$ 1 . Since $K_{\theta}=K_{z}\oplus zK_{z\frac{z-a}{1-\overline{a}z}},$ we take the basis $e_{1}(z)=1,\;e_{2}(z)=z/(1-\overline{a}z)$ and $e_{3}(z)=z^{2}/(1-\overline{a}z).$ Then it holds that
$A_{v}e_{1}=z=e_{2}-\overline{a}e_{3},\;A_{v}e_{2}=e_{3},\;A_{v}e_{3}=A_{v}(ae_{2}+\theta/z)=ae_{3}+ve_{1}.$
So $A_{v}$ is represented by the matrix
$[A_{v}]=\left(\begin{array}[]{ccc}0&0&v\\ 1&0&0\\ -\overline{a}&1&a\end{array}\right).$
Its eigenvalues satisfy $(\lambda-a)\lambda^{2}-v(1-\lambda\overline{a})=0$ or $\theta(\lambda)=v.$ And let $[B_{v}]=[A_{v}]^{*}$, we deduce that
$([B_{v}]-\overline{\lambda}I)k_{\lambda}^{\theta}$ $=\left(\begin{array}[]{ccc}-\overline{\lambda}&1&-{a}\\ 0&-\overline{\lambda}&1\\ \overline{v}&0&-\overline{\lambda}+\overline{a}\end{array}\right)\left(\begin{array}[]{c}1\\ \frac{\overline{\lambda}}{1-a\overline{\lambda}}\\ \frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}\end{array}\right)=\left(\begin{array}[]{c}-\overline{\lambda}+\frac{\overline{\lambda}}{1-a\overline{\lambda}}-\frac{a\overline{\lambda}^{2}}{1-a\overline{\lambda}}\vspace{1mm}\\ -\frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}+\frac{\overline{\lambda}^{2}}{1-a\overline{\lambda}}\vspace{1mm}\\ \overline{v}+\frac{\overline{\lambda}^{2}(\overline{a}-\overline{\lambda})}{1-a\overline{\lambda}}\end{array}\right)=0$
since $\theta(\lambda)=v.$ Thus $k_{\lambda}^{\theta}$ is an eigenvector of $B_{v}$ on $K_{\theta}$ with respect to the eigenvalue $\overline{\lambda}$ and then $Ck_{\lambda}^{\theta}$ is an eigenvector of $A_{v}$ on $K_{\theta}$ with respect to the eigenvalue $\lambda$.
(2)  Let $\theta(z)=z\left(\frac{z-a}{1-\overline{a}z}\right)^{2}$ when $10<|a|<$ 1 . Since $K_{\theta}=K_{z\frac{z-a}{1-\overline{a}z}}\oplus z\frac{z-a}{1-\overline{a}z}K_{\frac{z-a}{1-\overline{a}z}},$ the basis is $e_{1}(z)=1/(1-\overline{a}z),\;e_{2}(z)=z/(1-\overline{a}z)$ and $e_{3}(z)=z(z-a)/(1-\overline{a}z)^{2}.$ Then we deduce that
$A_{v}e_{1}=e_{2},\;A_{v}e_{2}=-\overline{a}ve_{1}+(a+(\overline{a})^{2}v)e_{2}+(1-|a|^{2})e_{3},\;A_{v}e_{3}=ve_{1}-\overline{a}ve_{2}+ae_{3}.$
That is, $A_{v}$ behaves as the following matrix
$[A_{v}]=\left(\begin{array}[]{ccc}0&-\overline{a}v&v\\ 1&a+(\overline{a})^{2}v&-\overline{a}v\\ 0&1-|a|^{2}&a\end{array}\right).$
Meanwhile, for the adjoint $[B_{v}]=[A_{v}]^{*}$, it holds that
$([B_{v}]-\overline{\lambda}I)k_{\lambda}^{\theta}$ $=\left(\begin{array}[]{ccc}-\overline{\lambda}&1&0\\ -a\overline{v}&-\overline{\lambda}+\overline{a}+a^{2}\overline{v}&1-|a|^{2}\\ \overline{v}&-a\overline{v}&-\overline{\lambda}+\overline{a}\end{array}\right)\left(\begin{array}[]{c}\frac{1}{1-a\overline{\lambda}}\\ \frac{\overline{\lambda}}{1-a\overline{\lambda}}\\ \frac{\overline{\lambda}(\overline{\lambda}-\overline{a})}{(1-a\overline{\lambda})^{2}}\end{array}\right)$
$=\left(\begin{array}[]{c}0\vspace{1mm}\\ -a\overline{v}+\overline{\lambda}\frac{\overline{a}-\overline{\lambda}}{1-a\overline{\lambda}}\frac{1-a\overline{\lambda}-1+|a|^{2}}{1-a\overline{\lambda}}\vspace{1mm}\\ \overline{v}-\overline{\lambda}\frac{(\overline{\lambda}-\overline{a})^{2}}{(1-a\overline{\lambda})^{2}}\end{array}\right)=\left(\begin{array}[]{c}0\vspace{1mm}\\ -a\overline{v}+a\overline{\lambda}\frac{(\overline{a}-\overline{\lambda})^{2}}{(1-a\overline{\lambda})^{2}}\vspace{1mm}\\ 0\end{array}\right)=0$
since $\theta(\lambda)=v.$ This also exhibits Proposition 3.5 and Remark 3.6.
Proposition 3.5, along with (1.2), gives the point spectrum and eigenvectors of $A_{z}^{\mathcal{M}}$ on $hK_{\theta}$.
Theorem 3.8.
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then $\sigma_{p}(A_{z}^{\mathcal{M}})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}$ with $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. For each $\lambda\in\sigma_{p}(A_{z}^{\mathcal{M}}),$ the corresponding eigenvector is $h(Ck_{\lambda}^{\theta})=h\widetilde{k_{\lambda}^{\theta}}.$
Considering $A_{v}$ as a one-dimensional perturbation of $S_{\theta}$, its spectrum must be the union of the non-Fredholm points of $S_{\theta}$ and its own eigenvalues, that is,
$\sigma(A_{v})=\sigma_{p}(A_{v})\cup\sigma_{e}(S_{\theta})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}\cup(\sigma(\theta)\cap\mathbb{T}),$
which is a special case of [21, Theorem 14.1] with the inner function vanishing at zero. Therefore, the preceding analysis enables us to derive a complete description for the spectrum of $A_{z}^{\mathcal{M}}$ on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
Theorem 3.9.
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then
$\sigma(A_{z}^{\mathcal{M}})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}\cup(\sigma(\theta)\cap\mathbb{T})$
where $v=\langle\theta,|h|^{2}\rangle$ with $|v|<1$.
4. Invariant subspaces of the compressed shifts on nearly invariant subspaces
In this section, we aim to characterize the invariant subspaces of our c.n.u. operator $A_{v}$ or $B_{v}$ with $|v|<1$ on a model space $K_{\theta}$ with $\theta(0)=0$. As an application, we completely solve Question 1 by describing all the invariant subspaces of the compressed shift $A_{z}^{\mathcal{M}}$ on arbitrary nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$.
For $f\in K_{\theta}$ with $\theta(0)=0,$ (2.6) demonstrates
$B_{v}f=S^{*}f+\overline{v}f(0)\theta/z.$
Write $B_{v}=S^{*}+F$ with $F(f)=\overline{v}f(0)\theta/z$ for short.
Suppose $\mathcal{N}\subseteq K_{\theta}$ is an invariant subspace for $B_{v}$, and that $f\in\mathcal{N}$ with $f(0)=0$. Then $S^{*}f=B_{v}f-Ff=B_{v}f\in\mathcal{N}.$ So $\mathcal{N}$ is a nearly $S^{*}$-invariant subspace and by Hitt’s theorem (Theorem 1.1) it can be written as $gK_{\phi}$ with unit norm $g\in\mathcal{N}$ orthogonal to $\mathcal{N}\cap zH^{2}(\mathbb{D})$ and $\phi$ is inner such that $\phi(0)=0.$ So we next explore when $gK_{\phi}\subseteq K_{\theta}$, and when it is invariant under $B_{v}$.
We note $gK_{\phi}$ is a nearly $S^{*}$-invariant subspace, then it is obvious that $g\in\mathcal{C}(K_{\phi}),$ which means that $|g|^{2}dm$ is a Carleson measure for $K_{\phi},$ that is, $gK_{\phi}\subseteq L^{2}(\mathbb{T}).$
Motivated by the above questions, we recall the result from [9] or the more general one in [2].
Theorem A [9, Theorem 3.1] For inner $\phi$ and $\theta$ and $g\in H^{2},$ the following are equivalent:
$(i)$ $gK_{\phi}\subseteq K_{\theta}$;
$(ii)$ $gS^{*}\phi\in K_{\theta},$ and $g\in\mathcal{C}(K_{\phi});$
$(iii)$ $g\in\ker T_{\overline{z}\overline{\theta}\phi}\cap\mathcal{C}(K_{\phi}).$
Thus Theorem A entails us to have $gK_{\phi}\subseteq K_{\theta}$ if and only if $gS^{*}\phi=g\phi/z\in K_{\theta},$ which, since $\phi(0)=0,$ is the same as $g\overline{z}\phi\overline{\theta}\in\overline{zH^{2}}$, or $\overline{g}\overline{\phi}\theta\in H^{2}.$
Next we consider $B_{v}(gK_{\phi})\subseteq gK_{\phi}$. For $k\in K_{\phi},$ it holds that
$B_{v}(gk)$ $=S^{*}(gk)+\overline{v}g(0)k(0)\theta/z$
$=\frac{g(z)k(z)-g(0)k(0)+\overline{v}g(0)k(0)\theta}{z}$
$=\frac{g(z)(k(z)-k(0))}{z}+\frac{(g(z)-g(0))k(0)}{z}+\frac{\overline{v}g(0)k(0)\theta}{z}$
$=\frac{g(z)(k(z)-k(0))}{z}+k(0)B_{v}g.$
The first term is always in $gK_{\phi},$ and without loss of generality we may take $k(0)=1$ (for if $k(0)=0$ there is nothing to prove). So a necessary and sufficient condition is $B_{v}g\in gK_{\phi}.$ In sum, the above steps present a characterization for invariant subspaces of $B_{v}$ on $K_{\theta}$ with $\theta(0)=0.$
Theorem 4.1.
Let $\mathcal{N}$ be an invariant subspace of $B_{v}$ with $|v|<1$ on $K_{\theta}$ with $\theta(0)=0$, then $\mathcal{N}$ is nearly $S^{*}$-invariant and then $\mathcal{N}=gK_{\phi}$ is an invariant subspace for $B_{v}$ if and only if $\theta\overline{g\phi}\in H^{2}$ and $(B_{v}g)/g\in K_{\phi}$.
The finite-dimensional invariant subspaces of $A_{v}$ or $B_{v}$ acting on a model space $K_{\theta}$ exhibit a simpler structure than others. This is particularly evident when the operator possesses at least one eigenvalue, as the corresponding eigenspace constitutes a nontrivial invariant subspace.
However, if $\theta$ is a Blaschke product of degree greater than $2$ and some eigenvalue has multiplicity $n>1$, then the generalized eigenspaces also serve as invariant subspaces. More precisely, suppose $\overline{\lambda}$ is an eigenvalue of $B_{v}$ with multiplicity $n>1$, then $\ker(B_{v}-\overline{\lambda})^{2}$ is an invariant subspace of $B_{v}$, which need not coincide with the eigenspace. Consequently, generalized eigenvectors are necessary to describe the invariant subspaces, particularly for high-order kernels such as $\ker(B_{v}-\overline{\lambda}I)^{n+1},\;n\geq 1.$ Analogous results hold for the operator $A_{v}$.
Now, for a function $f:\;\mathbb{C}\rightarrow\mathbb{C},$ let $Z(f)$ denote the zero set of $f$, i.e., the collection of all $z\in\mathbb{C}$ such that $f(z)=0$. The following theorem addresses the location of $Z(B^{\prime})$ for a Blaschke product (not necessarily finite) $B$.
Theorem B [3, Theorem 2.1] Let $B$ be a Blaschke product. Then the set $Z(B^{\prime})\cap\overline{\mathbb{D}}$ is included in the closed convex hull of $Z(B)\cup\{0\}.$
According to Theorem B, when $\theta$ is a Blaschke product of degree $n\geq 2,$ there exists $v\in\mathbb{D}$ such that the equation $\theta(\lambda)=v$ has a multiple root. Consequently, the eigenvalue $\overline{\lambda}$ of $B_{v}$ has algebraic multiplicity greater than $1$. For example, taking $\theta(z)=z\frac{z-\frac{1}{2}}{1-\frac{z}{2}}$ and $v=4\sqrt{3}-7,$ the equation $\theta(\lambda)-v=0$ has a double root $\lambda=2-\sqrt{3},$ which means $B_{v}$ on $K_{\theta}$ has an eigenvalue $2-\sqrt{3}$ with multiplicity 2. More generally, for a degree-2 Blaschke product $\theta$, the eigenvectors of $B_{v}$ corresponding to a double eigenvalue $\overline{\lambda}$ are precisely scalar multiples of $k_{\lambda}^{\theta}.$
Example 4.2.
Let $\theta(z)=z\frac{z-a}{1-\overline{a}z}$ when $10<|a|<$ 1 . Write $e_{1}(z)=1/(1-\overline{a}z)$ and $e_{2}(z)=z/(1-\overline{a}z).$ So $A_{v}e_{1}=e_{2}$ and $A_{v}e_{2}=ve_{1}+(a-v\overline{a})e_{2}.$ Therefore $A_{v}$ behaves as the matrix
$[A_{v}]=\left(\begin{array}[]{cc}0&v\\ 1&a-v\overline{a}\end{array}\right).$
The eigenvalues satisfy $\theta(\lambda)=v$ and $\sigma(A_{v})=\sigma_{p}(A_{v})=\{\lambda\in\mathbb{D}:\;\theta(\lambda)=v\}.$ Suppose $x_{1}e_{1}+x_{2}e_{2}$ is the eigenvector of $B_{v}$ with respect to $\overline{\lambda}$ and then it holds that
$\left(\begin{array}[]{cc}\overline{\lambda}&-1\\ -\overline{v}&\overline{\lambda}-\overline{a}+\overline{v}a\end{array}\right)\left(\begin{array}[]{c}x_{1}\\ x_{2}\end{array}\right)=0.$
Solving the equations, it yields that the eigenvector $x_{1}e_{1}+x_{2}e_{2}=tk_{\lambda}^{\theta}$ with $t\in\mathbb{C}$.
Going beyond Example 4.2, let us consider the case where the eigenvalue $\overline{\lambda}$ of $B_{v}$ has multiplicity $n+1$ for some $n\geq 2$, with $\theta(\lambda)=v$. And then $\theta^{(k)}(\lambda)=0$ for $k=1,\cdots,n.$ We need to formulate the $n$-th derivative kernel $(k_{\lambda}^{\theta})^{(n)}$ on $K_{\theta}$ by the $n$-th derivative kernel in $H^{2}$ as below
$k_{\lambda}^{(n)}(z)=\frac{n!z^{n}}{(1-\overline{\lambda}z)^{n+1}}$
satisfying $\langle f,k_{\lambda}^{(n)}\rangle=f^{(n)}(\lambda)$ for all $f\in H^{2}.$ And then
$(k_{\lambda}^{\theta})^{(n)}(z)$ $=\langle k_{\lambda}^{(n)}(w),k_{z}^{\theta}(w)\rangle=\langle\frac{n!w^{n}}{(1-\overline{\lambda}w)^{n+1}},\frac{1-\overline{\theta(z)}\theta(w)}{1-\overline{z}w}\rangle$
$=\frac{n!z^{n}}{(1-\overline{\lambda}z)^{n+1}}-\theta(z)\langle\frac{n!w^{n}}{(1-\overline{\lambda}w)^{n+1}},\frac{\theta(w)}{1-\overline{z}w}\rangle$
$=\frac{n!z^{n}}{(1-\overline{\lambda}z)^{n+1}}-\theta(z)\sum_{k=0}^{n}C_{n}^{k}\frac{k!z^{k}}{(1-\overline{\lambda}z)^{k+1}}\overline{\theta^{(n-k)}(\lambda)}.$
Since $\theta(\lambda)=v$ and $\theta^{(k)}(\lambda)=0$ for $k=1,\cdots,n,$ it yields that
$(k_{\lambda}^{\theta})^{(n)}(z)=\frac{n!z^{n}(1-\overline{v}\theta)}{(1-\overline{\lambda}z)^{n+1}},$ (4.1)
ensuring $\langle f,(k_{\lambda}^{\theta})^{(n)}\rangle=f^{(n)}(\lambda)$ for all $f\in K_{\theta}.$ At this time, the action of $i$-th iteration of $B_{v}$ is
$B_{v}^{i}(k_{\lambda}^{\theta})^{(n)}(z)=\frac{n!z^{n-i}(1-\overline{v}\theta)}{(1-\overline{\lambda}z)^{n+1}},\;\;i=1,2,\cdots,n.$
And the action of $n+1$-st iteration of $B_{v}$ is
$B_{v}^{n+1}(k_{\lambda}^{\theta})^{(n)}(z)$ $=n!S^{*}\left(\frac{1-\overline{v}\theta}{(1-\overline{\lambda}z)^{n+1}}\right)+\overline{v}\theta/z$
$=n!\frac{(1-\overline{v}\theta)(1-(1-\overline{\lambda}z)^{n+1})}{z(1-\overline{\lambda}z)^{n+1}}.$
Based on the above calculations, we deduce that
$(B_{v}-\overline{\lambda}I)^{n+1}(k_{\lambda}^{\theta})^{(n)}(z)$ $=\sum_{i=0}^{n+1}C_{n+1}^{i}(-\overline{\lambda})^{n+1-i}B_{v}^{i}(k_{\lambda}^{\theta})^{(n)}(z)$
$=n!\sum_{i=0}^{n}C_{n+1}^{i}(-\overline{\lambda})^{n+1-i}\frac{z^{n-i}(1-\overline{v}\theta)}{(1-\overline{\lambda}z)^{n+1}}+n!\frac{(1-\overline{v}\theta)(1-(1-\overline{\lambda}z)^{n+1})}{z(1-\overline{\lambda}z)^{n+1}}$
$=n!\frac{1-\overline{v}\theta}{(1-\overline{\lambda}z)^{n+1}}\left(-\overline{\lambda}\sum_{i=0}^{n}C_{n+1}^{i}(-\overline{\lambda}z)^{n-i}+(1-(1-\overline{\lambda}z)^{n+1})/z\right)$
$=n!\frac{1-\overline{v}\theta}{(1-\overline{\lambda}z)^{n+1}}\left(-\overline{\lambda}\sum_{i=0}^{n}C_{n+1}^{i}(-\overline{\lambda}z)^{n-i}+\overline{\lambda}\sum_{i=0}^{n}C_{n+1}^{i}(-\overline{\lambda}z)^{n-i}\right)=0.$
This implies the $n$-th derivative kernel $(k_{\lambda}^{\theta})^{(n)}$ for $\lambda$ in (4.1) is a generalized eigenvector of $B_{v}$ when the eigenvalue $\overline{\lambda}$ has multiplicity $n+1$. That is,
Proposition 4.3.
Let $B_{v}:=A_{|h|^{2}\bar{z}}^{\theta}$ be the adjoint truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$. Suppose the eigenvalue $\overline{\lambda}$ of $B_{v}$ has multiplicity $n+1$ for $n\geq 2$ and satisfies $\theta(\lambda)=v,$ then $(k_{\lambda}^{\theta})^{(n)}=\frac{n!z^{n}(1-\overline{v}\theta)}{(1-\overline{\lambda}z)^{n+1}}$ is a generalized eigenvector of $B_{v}$.
At this stage, we encounter another key question:
Question 2: How can we characterize the nontrivial invariant subspaces for $A_{v}$ or $B_{v}$ on $K_{\theta}$ with $\theta(0)=0$, particularly when $\theta(\lambda)$ never takes the value $v\in\mathbb{D}$?
As an example, consider the singular inner function $u(z)=\exp(\frac{z-1}{z+1})$, which satisfies $u(0)=1/e$ and never vanishes. Defining the Frostman shift of the inner function $u$ as
$\theta(z):=\frac{u(z)-1/e}{1-u(z)/e}=u_{\frac{1}{e}}(z),$
we observe $\theta$ is inner, $\theta(0)=0$, and $\theta$ never takes the value $v:=-1/e$. If the spectrum of $A_{v}$ or $B_{v}$ is disconnected then it always has invariant subspaces. But the spectrum in this example is one point on the circle, so how do we answer Question 2?
At this moment, the unitary equivalence in the diagram (3.5), together with, the fact the invariant subspaces of $S_{\theta_{v}}$ in $K_{\theta_{v}}$ are precisely $K_{\theta_{v}}\ominus K_{\phi}$, where $\phi$ divides $\theta_{v}$ (see, e.g.[11, Proposition 9.14]) establish the invariant subspaces for $A_{v}$ as below.
Theorem 4.4.
Let $A_{v}:=A_{|h|^{2}z}^{\theta}$ be the truncated Toeplitz operator on a model space $K_{\theta}$ with inner function $\theta$, $\theta(0)=0$ and $v=\langle\theta,|h|^{2}\rangle$ satisfying $|v|<1$, then the invariant subspaces of $A_{v}$ in $K_{\theta}$ are precisely $J_{v}^{-1}(K_{\theta_{v}}\ominus K_{\phi})$, where $\phi$ divides $\theta_{v}$ and $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is defined in (3.4).
Finally, using (1.2) we can completely answer Questions 1 and 2 with the following theorem.
Theorem 4.5.
Let $A_{z}^{\mathcal{M}}$ be the compressed shift on a nearly $S^{*}$-invariant subspace $\mathcal{M}=hK_{\theta}$, where $h$ is an extremal function and $\theta$ is the associated inner function satisfying $\theta(0)=0.$ Then every $A_{z}^{\mathcal{M}}$-invariant subspace of $\mathcal{M}$ is of the form $hJ_{v}^{-1}(K_{\theta_{v}}\ominus K_{\phi})$, where $v=\langle\theta,|h|^{2}\rangle$ with $|v|<1$, $\phi$ is an inner function dividing the Frostman shift $\theta_{v}$ of $\theta$, and $J_{v}^{-1}:\;K_{\theta_{v}}\rightarrow K_{\theta}$ is the inverse of the Crofoot transform.
To conclude this paper, we present a concrete example to illustrate our main results.
Example 4.6.
Consider the nearly $S^{*}$-invariant subspace $\mathcal{M}=(1+z)K_{z^{2}}$. The unnormalized reproducing kernel at 0, denoted by $k(z):=P_{\mathcal{M}}1$, can be explicitly determined by expressing it in the form $k(z)=(z+1)(p+qz)$ and doing evaluations at 0 so that
$\langle(z+1)(p+qz),z+1\rangle=1\quad\text{and}\quad\langle(z+1)(p+qz),z(z+1)\rangle=0.$
This yields that $p=2/3$ and $q=-1/3$. Now $\|k\|^{2}=\langle k,k\rangle=k(0)=2/3$, leading to the normalized reproducing kernel
$h(z)=\sqrt{\frac{3}{2}}\frac{(2-z)(z+1)}{3}=\frac{(2-z)(z+1)}{\sqrt{6}}.$ (4.2)
Now the model space $K_{\theta}$ must contain $1/(2-z)$ and $z/(2-z)$, which enables us to choose
$\theta(z)=z\frac{z-1/2}{1-z/2}.$ (4.3)
So $(1+z)K_{z^{2}}=hK_{\theta}$ with extremal function $h$ in (4.2) and associated inner function $\theta$ in (4.3).
We then compute
$v=\langle\theta,|h|^{2}\rangle=\langle h\theta,h\rangle=\frac{1}{6}\left(2(z+1)(z-1/2)z,(2-z)(z+1)\right)=-\frac{1}{3}.$
Solving the quadratic equation $\theta(z)=v$ yields the eigenvalues
$a=\frac{1+i\sqrt{2}}{3},\;\;b=\frac{1-i\sqrt{2}}{3}.$
By Theorem 3.9, the spectrum of $A_{z}^{\mathcal{M}}$ is
$\sigma(A_{z}^{\mathcal{M}})=\sigma_{p}(A_{z}^{\mathcal{M}})=\{a,b\},$
and Theorem 3.8 ensures the functions $(z-a)(z+1)$ and $(z-b)(z+1)$ are the eigenvectors corresponding to the eigenvalues $b$ and $a$, respectively.
The Frostman shift $\theta_{v}$ of $\theta$, as defined in (3.3), is given by
$\theta_{v}(z)=\frac{\theta(z)-v}{1-\overline{v}\theta(z)}=\frac{3z^{2}-2z+1}{z^{2}-2z+3}=\frac{z-a}{1-bz}\frac{z-b}{1-az}:=\theta_{1}\theta_{2},$
where
$\theta_{1}(z)=\frac{z-a}{1-bz},\;\;\theta_{2}(z)=\frac{z-b}{1-az}.\;\;$
Furthermore, the model space $K_{\theta_{v}}$ admits two orthogonal decompositions
$K_{\theta_{v}}=K_{\theta_{1}}\oplus\theta_{1}K_{\theta_{2}}=K_{\theta_{2}}\oplus\theta_{2}K_{\theta_{1}}.$
Theorem 4.5 implies the only nontrivial invariant subspaces of $A_{z}^{\mathcal{M}}$ on $hK_{\theta}$ are $hJ_{v}^{-1}(\theta_{1}K_{\theta_{2}})$ and $hJ_{v}^{-1}(\theta_{2}K_{\theta_{1}})$, where $J_{v}^{-1}$ is defined in (3.4). Noting that
$\theta_{1}K_{\theta_{2}}=\mathbb{C}\frac{\theta_{1}}{1-az}\;\;\mbox{and}\;\;\theta_{2}K_{\theta_{1}}=\mathbb{C}\frac{\theta_{2}}{1-bz},$
we explicitly calculate
$J_{v}^{-1}\left(\frac{\theta_{1}}{1-az}\right)$ $=\frac{2\sqrt{2}}{3}\frac{\theta_{1}}{1-\frac{1}{3}\theta_{v}}\frac{1}{1-az}=\frac{3\sqrt{2}}{2}\frac{z-a}{2-z},$
$J_{v}^{-1}\left(\frac{\theta_{2}}{1-bz}\right)$ $=\frac{2\sqrt{2}}{3}\frac{\theta_{2}}{1-\frac{1}{3}\theta_{v}}\frac{1}{1-bz}=\frac{3\sqrt{2}}{2}\frac{z-b}{2-z}.$
Consequently, the nontrivial invariant subspaces of $A_{z}^{\mathcal{M}}$ are
$hJ_{v}^{-1}(\theta_{1}K_{\theta_{2}})=\mathbb{C}(1+z)(z-a)\;\;\mbox{and}\;\;hJ_{v}^{-1}(\theta_{2}K_{\theta_{1}})=\mathbb{C}(1+z)(z-b),$
which are precisely the eigenspaces corresponding to the eigenvalues $b$ and $a$, respectively.
Acknowledgements   The work was supported in part by the National Natural Science Foundation of China (Grant No. 12471126).